Running Log

20/10/2016  
1215 Depart home by road to Littlehampton.
1600 Following road closures and detours, arrive Premier in Littlehampton.
1700 Contacted D McAndie and Rib Charter wrt departure times for tomorrow, 1100  departure confirmed.
21/10/2016  
0900 Check out from hotel and proceed to Shoreham. Park at Shoreham Yacht Club as instructed by RIB Charter.
1100 Depart on CTV Galaxy to Stemat Spirit
1130 Arrive and board Stemat Spirit, vessel is ploughing and burying cable on arrival. D McAndie has issued C of A for that operation. Weather good and forecast to remain so.
1800 Evening Meal, vessel has been ploughing steadily throughout the day on an anchor spread.
22/10/2016  
0650 At KP 2.65
0730 Receive vessel induction.
1000 Morning meeting, vessel expects to be complete by mid-week 44. Cut point is at KP12, steady progress at around 100m/he being made. Weather good and expected to remain so.
2150 Plough Layback now 72 metres.
23/10/2016  
0630 KP 4.625
1000 Morning meeting, steady progress being made, cut predicted for Wednesday/Thursday, v/l in Rotterdam by mid/end week 44.
24/10/2016  
0700 Previous 24 hours cable laid 1948m (to midnight)
1000 Steady progress being made, ETA Rotterdam unchanged. Weather outlook good.
25/10/2016  
0630 At KP 8.874, 1277m laid in previous 24 hours (to midnight)
1000 Morning meeting, steady progress being made and no change to existing ETA Holland. Internet down, technicians working on the problem.

Weather outlook good. Ploughing on anchor spread all day.

26/10/2016  
0620 At KP 11.061
0930 Inspect supplied abandonment kits and long-term seal arrangements. The abandonment/pull in heads are very ‘robustly’ engineered. Some concerns that they may be difficult to fit to the cable. Four kits available onboard.
1000 Meeting, cutting later today. All preparations in hand. Weather outlook good. Vessel will swap to DP operations following cut.
1713 Vessel readying to recover plough.
1823 Crane wire connected to plough, taking weight.
1905 Plough on deck
2024 Vessel in position for cutting cable.
2055 Commence cutting cable
2125 Cutting complete (777m section to be  removed)
2230 Chinese fingers applied to both ends of cut cable
2320 Cable recovered sufficiently to stopper off and carry out sealing procedure.
27/10/2016  
0030 Cable testers complete testing from within carousel
0100 Testing of cable to shore end completed.
0420 OTDR testing complete
1000 Daily meeting, extra cable sealers are now available to assist the one embarked sealer.
1530 Sealing complete commence fitting abandonment cap.
1545 Head completed
1650 Commence laying (good) cable end down on seabed.
1710 Head overboarded.
1750 Cable lay down anchor overboarded.
1840 Commence sealing faulty cable end.
2340 V/L in DP mode, anchors demobilised to anchor handlers for wet store.
28/10/2016  
0045 Sealing complete
0220 Vessel in position to start laydown of faulty cable
0250 Overboarding laydown anchor.
0350 Cable connected to laydown rigging,
0700 Vessel in 2nd cut positon, cable marked by client at cut point.
0800 Testing cables from carousel
0905 Tests complete preparing to cut cable.
0915 To 0945, cutting cable.
1000 Commence sealing faulty cable. Morning meeting, weather forecast to remain good.
1240 Commence fitting Chinese finger to faulty cable 2nd end.
1550 Following the detection of further faults, EON instruct that a further 100m of cable to be cut out, this will be done in 10m lengths.
1555 Commence cutting.
1920 Complete cutting
2050 Testers report remaining cable as good. Commence sealing good end of remaining cable
29/10/2016  
0250 Sealing complete
0450 Cable threaded through plough, side-lacing in place.
0510 Cable (good) on chute
0520 Freelaying to RPL
0710 Vessel at plough launch position.
0810 Plough lifted clear of deck
0825 Plough in splash zone.
0840 Plough unlatched from crane.
1000 Morning meeting, nothing of note, weather remains good.
1040 Ploughing away.
30/10/2016  
0725 KP13942,
1000 Morning meeting, vessel expecting to continue ploughing throughout the day. Did not retard clocks on hour to GMT. Vessel now one ahead of UK time. Weather remaining good.
31/10/2016  
0500 All stopped at plough recovery position
0900 Plough connected to crane wire.
0950 Plough on deck
1420 Following cable testing and free lay of final length of cable, commence cutting cable.
1440 Complete cable cut.
1525 Chinese fingers fitted. Cable testing complete.
1700 Commence sealing cable 2nd end.
01/11/2016  
0125 Long-term seal and abandonment cap fitted.
0140 Laydown pennant connected.
0142 Overboarding
0240 2nd end on bottom.
0250 Vessel proceeds to recover damaged section.
1210 Good/bad ends recovered to deck. Good cable stoppered off.
1835 Longterm seal and abandonment head fitted to good end.
1905 Overboarding good end connected to laydown rigging.
2005 Anchor overboard
2140 Depart vessel on crewboat Galaxy.
2200 (GMT) Check into Best Western at Hove
02/11/2016  
0830 Check out of hotel, commence road passage to Margate
1150 Arrive home.
   
   
  Export Cable #Two Attendance
10/11/16  
0420 Depart home by taxi to LCY
0745 Depart on Cityjet Flight to Amsterdam
1130 Board Stemat Spirit at Amerikahaven
1300 Attended HIRA for installation of 2nd cable
1900 Cable testing completed, testers depart vessel. Faults identified. EON aware and require the cable to be installed without cutting out any sections.
11/11/16  
0600 Vessel remains in port.
1000 Morning meeting sailing said to be 1500 today.
1350 Transport to immigration office and back to vessel. Instructions from Rampion to delay sailing due to contractual concerns wrt integrity of the cable.
1700 Decision taken to sail in order to use forecast weather window.
2115 Sailaway C of A issued C13725.
2250 Departed Amsterdam
12/11/16  
0200 Clocks retarded 1 hour to GMT
1000 Morning meeting, weather window now looking marginal on Wednesday. Crane reported as operational but two technicians will be boarding tomorrow to carry out maintenance. Voyage to location continues. ETA location 0600.
1810 Boat drill.
2230 MWS to bed.
13/11/16  
0900 On location in DP weather very good.
1000 Morning meeting, weather window gone. TMS to be installed to Nova K and Yvonne W and checked out on Sea Bravo. DP Trials will be conducted after that. Spring tides at play and a ‘supermoon’ on Monday.
14/11/16  
0145 Called for the start of DP Field arrival trials
0445 Complete DP trials, no issues noted.
1000 Morning meeting, no weather window in sight. OCM leaving for a meeting in Holland tomorrow, beach pull in teams may be stood down. Worsening weather at the end of the week may require the AHTs to be sent to shelter and indeed Stemat Spirit may seek shelter in the Solent.
1200 Main crane incident (tbc) control issues resulting in main hoist heaving out of control and damaging sheave housing and auto stop plate.
1600 Receive information from Rampion that vessel will carry out repairs in Southampton and also it is suggested, wait there for an improvement in weather prospects. This has not been confirmed by VBMS however.
2230 MWS to bed, vessel remains at location on DP with tugs at anchor nearby in attendance.
15/11/16  
0600 MWS up and about, vessel remains on location in DP with tugs in attendance at anchor. Weather very good, seas <0.5m.
1000 Morning meeting.

As a result of the crane incident yesterday, two manufacturer’s technicians and the head of Stemat’s technical department will join the vessel this afternoon to investigate the incident and its cause. One aboard the vessel will sail to Southampton to conduct repairs and any re-certification that is required. The weather forecast is very poor for the weekend so the vessel will remain in port sheltering until the weather improves.

Spare cable abandonment kits will be delivered during the port-call.

Client rep requested that VBMS produce contingency plans to free lay the cable beyond pit three following shore pull in. A MOC will be produced.

1835 Informed that the vessel will proceed to Rotterdam for crane repairs.
2115 Vessel departs location to Rotterdam
16/11/16  
0200 Clocks advanced 1 hour to Dutch time (GMT + 1)
0600 MWS up and about, vessel proceeding at 7 knots to Rotterdam. ETA pilotage station midnight.
0800 Walk the deck and inspect carousel and cable path, all well.
1000 Morning meeting, the crane incident investigation continues. There are two Stemat technicians and two Heila technicians onboard working on the investigation. Work continues 24 hours a day.

A root cause is sought so the investigation may take some time. That said, VBMS are hoping to sail from Rotterdam by Sunday this weekend and return to location following crane repairs and re-certification.

No weather window permitting pull-in operations to commence has been identified yet.

1700 Walk the deck and inspect carousel and cable path, all well.
2330 MWS to bed
17/11/16  
0400 Berth at

Waal / Eemhaven, Rotterdam
Harbournumber: 2732
Address: Den Hamweg 30, 3089 KK, Rotterdam

0730 Crane investigation kick off meeting starts. MWS and client not present. Around 14 people present by the look of it.
0900 Walk the deck, cable OK.
1000 Morning meeting. Crane investigation continues, Heila technicians in attendance.
  Vessel expects to remain at current berth until departure. Weather conditions on Sunday are forecast to be high and the port may be closed to navigation.

Weather outlook at Rampion is still not conducive to operations, no window identified.

1200 Walk the deck, crane investigation ongoing.
2300 MWS to bed.
18/11/16  
0640 MWS up and about.
0730 Walk the deck, all well, crane investigations continue. Weather dam and breezy, temp circa 3 degrees.
1000 Morning meeting: Crane investigation and repair preparations continue. Main lift wire will be cropped and re-socketed. Sheave will be polished, sheave guides will be renewed.

Load-testing expected to under-taken on Monday, possibly Tuesday. GM will be sending a crane specialist to attend testing.

No weather window identified at Rampion as yet. There is some variation between the Stormgeo and Meteo forecasts.

2300 MWS to bed
19/11/16  
0700 MWS up and about. News from the crane investigation. An error in the wiring of transformers in the crane has been identified as the cause of the uncontrolled ‘heave’. The incident has been repeated and the wiring fault has been confirmed as the root cause.
0900 Contractors on site to conduct re-socketing and sheave repairs.
1000 Morning meeting. Repairs are expected to be completed today. Diagnostics will continue on Sunday and load-tests on Monday. Hopefully a Monday night departure.
1230 Crane repairs complete.
1800 Walk decks, all well.
2230 MWS to bed
20/11/16  
0700 MWS up and about.
0800 Walk the decks, all well, very strong wind conditions.
1000 Morning meeting, crane diagnostics and function tests will continue through the day. Tomorrow load-tests will be conducted by Heila, overseen by Chief Engineer, reports will be sent to BV and in due course certificates issued.

A possible weather window is identified by Meteo for next week. Stormgeo do not.

1800 Walk the deck
   
2300 MWS to bed
21/11/16  
0630 MWS up and about
0700 Walk the deck, all well. Crane team not yet onboard.
0850 GM crane man Olav Fagerbakke boards, introduced to ship’s management team.
1000 Morning meeting. Load tests will be conducted today, departure still planned for this evening all being well. Load-test personnel and equipment have arrived at vessel
1120 Water bags attached to crane main hoist and being filled with water. Rigging certs checked, load cell calibration certificate sighted and in order.
1210 With 50.5 tonnes showing on load-cell overload tested, not successful on first attempt.

Subsequent tests of the overload (auto put down of a load that is determined to be in excess of cranes limits) appear to be satisfactory although some ‘delays’ in the crane’s reaction are noted.

1600 VBMS cable tester on board. Due to loss of data when downloading previous tests, they are going to repeat the TDR tests of the cable cores.
1620 GM crane specialist states that in ‘manual’ mode the crane is acceptable to launch and recover the plough for cable operations.
1720 GM crane specialist Olav Fagebakke reports that crane is working correctly on the main lift in auto mode. A report will follow in due course.
1850 Aux lift tested and functioning correctly in auto mode.
1900 Olav Fagebakke departs vessel.
1905 TD confirmed as 2200 tonight.
2205 Depart berth, outwards.
22/11/16  
0700 MWS up and about, vessel proceeding into stiff seas progress down to <2knots at times. ETA Rampion afternoon of the 23rd.
1000 Morning meeting, weather is easing gradually and the vessel will speed up as the effect of the head seas diminishes. Awaiting crane report from Heila.
2300 Vessel passage continues through the day.
23/11/16  
0200 Clocks retarded one hour to GMT
0700 MWS up and about, vessel off Eastbourne ETA Rampion 1200.
0910 Discussion with OCM, wrt weather, it’s a marginal call, still. VBMS have asked Meteo to adjust the forecast position of ‘Cable Landing’ from the current beach location to the first pit. Once done that forecast will be revisited with regards to the go no go decision making process.
1000 Morning meeting, the Meteo Cable Landing weather forecast has had its position changed to 50 deg 48’ North and 000 degrees 19’West, the position of the first pit. Both Stormgeo and Meteo forecasts show that conditions are out of limits to start operations.
1030 On location, tugs in attendance. Starboard boat lading being readied.
1235 Commence DP arrival trials. Sight crane socket certification.
1455 Complete DP arrival trials, MWS present throughout. No issues noted, all appears satisfactory. Sight toolbox talk summary (AHT pre-work briefing) for assisting vessels.
1620 Justin Hewlett boards.
1635 MOC received FB000042-VBMS-GEN-MOC-0012-5107_Revised weather conditions for pull-in, dated 12/10/2016. Send to project team in Aberdeen.
2300 MWS to bed.
24/11/16  
0700 MWS up and about, very windy from the North east. Walk the decks, cable and lay equipment appears in order.
1000 Morning meeting, weather window possible at the weekend, will monitor and react as required. Plough depressor requires calibrating. Crane report awaited. Request that the vessel forwards test documentation to GM Norway as agreed during the testing. Plough wear plate will be strengthened (additional plate added) when conditions allow the seafastening to be removed.
1600 Walk the deck wind still strong but all well.
2300 MWS to bed.
25/11/16  
0600 MWS up and about.
0700 Walk the deck, wind has eased, sea conditions still robust.
1000 Morning meeting, although a weather window is opening on Saturday morning it isn’t long enough for operations to get underway. The situation is under review and forecasts are carefully monitored. AHT Novak has a draft of 2.6m and Yvonne W a draft of 3.2m, their under-keel clearance is part of the decision making process.
1105 Until 1135, the kick off meeting/project overview, conducted initially in Amsterdam on November was for all intents and purposes, repeated. It seemed to take the same amount of time as the original meeting too.
1600 Receive by email a copy of the technical analysis (water depths and environmental conditions required compared to forecast conditions) as to why we are unable to start operations.
2300 MWS to bed
26/11/16  
0630 MWS up and about. Conditions have eased considerable overnight, wind now 10 knots from the NE, seas <1m
O850 Port boat landing installed. Plough depressor calibrated.
1000 Morning meeting. WOW still. Tuesday looking promising. Abandonment heads are in transit to vessel, probably arriving early in the week. Plough shear thickening plate will be welded today following locating plough to launch position on deck and re-seafastening.
1630 Walk the deck, all well.
2300 MWS to bed
27/11/16  
0700 MWS up and about. WOW
1000 Morning meeting. Preparations continue, anchors being wet-stored (from decks of AHTS), Spirit’s port anchor will be removed from anchor wire today in readiness for operations. Current forecasts indicate that vessel is likely be moored in the 2nd pit by 2100 on Tuesday evening.

Beacons being tested, plough welding continues. Crane report from VMS received.

1100 Carousel seafastening removed
1555 Cable being routed through plough.
1635 Cable head on chute, preparing to install side lace stocking.. Current sensor overboard and transmitting correctly.
1755 Side lace stocking being installed to cable shore end, immediately aft of the Chinese finger.
1930 Side-lace stocking completed.
28/11/16  
0700 MWS up and about
0800 Walk the deck, all OK.
1000 Morning meeting, weather looking very good for a commencement of ops on Tuesday evening, moving into pit #2 at 2100 and running anchors. Pull in ops to start the following day. Abandonment kits will come out with the cable jointers. There will be a total of three kits on board at that point. Cable testers will board at KP4.
2300 MWS to bed
29/11/16  
0700 MWS up and about. Wind has eased considerably and sea conditions also much improved.
0830 Walk decks. All well, preparations completed, still WOW
1000 Morning meeting, ops will commence this evening. Medium term forecast very promising. ECE personnel will join tomorrow.
1500 Stemat Spirit slowly manoeuvring along lay route towards KP4 holding point.
1600 At KP4, standing by.
1800 Slowly moving in to pit #2.
2105 In pit#2
2120 Commence running anchors, Nova K and Sea Bravo in attendance.
2127 #1 anchor on the bottom.
2137 Running #2
2220 Running #5 and #6 anchors.
2250 Running#3 anchor
2330 #7 on the bottom (in wet stow)
2333 Running #4
30/11/16  
0140 #7 anchor on bottom on the pull line. Tidal window for next operations 0840 to 1425.
0850 Transiting to #1 pit.
0950 In Pit #1, messenger passed to CTV Galaxy
0954 Messenger connected to riser, heaving in.
1020 Pipe at chute
1120 Pipe head at top of chute and stoppered off.
1252 Flange removed.
1255 C of A issued. C 13757 commencement of shore-end pull in of 2nd cable.
1307 Pull in wire connected to cable rigging.
1313 Shore pulling, weight on cable.
1318 Cable head enters pipe.
1334 Forty metres pulled through, deck crew commence greasing cable.
1632 Stop
1640 Pull completed.
1700 End cap being fitted to pipe.
1935 End cap secured, keeper plates in place, issue Plough Launch/initiation C of A 13758.
1945 Pipe stoppers removed.
2006 Plough clear of deck.
2110 Commence laying back to 65m
2125 Issue Lay and Burial C of A C13759. Weather conditions excellent, wind zero, sea flat.
2145 At 65m layback.
2150 Commence laying away towards Pit #2 (1m/min)
01/12/16  
0228 In Pit #2, MWS to bed
0715 MWS up and about.
0840 Vessel commences laying away from Pit #2
1030 Resetting anchors.
1055 CTV Galaxy alongside, brings two VBMS VIPs.
1105 Resetting anchors complete, tensioning up.
1106 Laying away.
1433 In pit #3 (Client reps changed out in the early afternoon, Justin Hewlett replaced by John Elliot)
1933 Tidal window opens, commence laying away from Pit #3
2254 Stop, need to reset Survey software.
2258 Continue laying away.
02/12/16  
0615 KP 3.167, MWS up and about, proceeding well at 3m/min
1000 Morning meeting, vessel will change over to DP mode between 0001 and 0600 tomorrow morning once he vessel is in deep enough water to allow 2m UKC (measured from bottom of azimuth thrusters). Weather forecast to build on Saturday afternoon and last through Sunday. Anchor handling tug Sea Brave will be replaced by AHT Viking on Sunday.
1230 KP 4 point approaching, watching cable for evidence of damage (in vicinity of factory mark 13km).  A ‘twist’ was detected in the cable about 50m offshore side from the 13km mark.
1620 Fibre optic testers from shore report that results (OTDR) show that all but two of the optic fibres are good. This contrary to earlier tests.
2300 MWS to bed, cable burial continues well.
03/12/16  
0010 Until 0230, convert to DP mode, all anchors removed apart from #7.
0650 KP 6.01, making good progress. Vessel now in DP mode in readiness for the worsening weather conditions forecast for later today. Burial achieved > 1.5metres.
0900 Decision taken to disconnect #7 anchor and WOW
1000 Morning meeting; will WOW until the morning of the 4th when conditions are forecast to improve.
1045 Altering heading to port, shaping up to the weather (Easterly),
2230 WOW MWS to bed.
04/12/16  
0500 Conditions easing vessel altering heading and moving back to lay-line.
0720 Commencing laying away.
0930 KP 6.7.
1000 Morning meeting, vessel currently laying at around 4 metres per minute which means, if rate of progress is maintained, that plough will be covered on Tuesday afternoon. Poor weather is forecast on Thursday.

The main AHT Sea Bravo will be replaced by AHT Viking this afternoon. Sea Bravo will depart site once the Viking’s Master has received a pre-ops tool box talk and the TMS (survey pack) is proved operational.

Sebastion Vassile, the VBMS cable tester will join today and reset the reference value on the cable OTDR monitoring system.

1400 AHT Viking takes over from Sea Bravo